How to solve 2x+3=-19
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

2x+3=-19
2x=-22
x=-11

Answer by checkley77(12844)

2x+3=-19
2X=-19-3
2X=-22
X=--22/2
X=-11 ANS.